Determine the pH of a 0.35 M aqueous solution of butylamine at 25.0 °C. The pKb of butylamine is 3.22.
1
Expert's answer
2019-10-21T07:16:29-0400

C( butylamine) = 0.35M= Cb

pKb=3.22 Kb=10^(-3.22)=0.0006

[OH-]= (Kb*Cb)^(1/2)=0.01449

pOH = -lg[0.01449] = 1.84

pH = 14 - pOH = 12.16